Jones Company has the following standards for its single product:

standard quantity standard price

direct materials 13 pounds per unit    $5.10 per pound
direct labor 6 hours per unit $16.00 per hour
variable overhead 6 hours per unit    $11.00 per hour

Jones Company reported the following information for the month of October:

1. 4,280 units were produced.
2. 63,000 pounds of direct materials were purchased at cost of $5.40 per
pound.
3. 27,200 direct labor hours were worked.
4. The actual variable overhead cost amounted to $303,800.
5. The actual direct labor cost amounted to $414,800.
6. There were no inventories of any type at October 1, however, the direct
materials inventory at October 31 totaled 6,000 pounds.

Calculate the direct labor rate variance for October. If the variance is
favorable, place a minus sign in front of your answer (i.e., -5000). If
the variance is unfavorable, enter your answer as a number (i.e., 5000).
Answer #1

Labor rate variance

= (SR - AR) * AH

= (16 - 414,800/27,200) * 27,200

= (16 - 15.25) * 27,200

= - 20,400 Favourable (don't forget the minus sign ;))